# Spin-up and spin distribution of stellar black holes grown by gas accretion in proto-stellar clusters

### 19 Mar 2026 v1

@@ 5 Conclusion
We have calculated the BH spin distribution generated by accretion in gaseous proto-stellar clusters within $\sim 10\,{\rm Myr}$ as gas is exponentially depleted by stellar feedback. We have mainly considered parameter values consistent with JWST observations of five proto-stellar clusters in the Cosmic Gems arc galaxy (Adamo et al., 2024) and also explored a broader parameter space.
Our accretion model incorporates the repeated disruption of BH accretion disks by gravitational perturbations in dense cluster environments. The disk orientation is determined by the acquired BH’s azimuthal velocity, and therefore each disruption reorients the disk. Our major results are summarized in Tables 1, 2, in Figure 5(b), and Eq. (31).
This equation describes a BH spin distribution as a function of BH mass with a remarkably consistent functional form; it persists for all cluster parameter values investigated here, with representative ranges given in Table 1. This behavior, depicted graphically in Figure 5(b), reveals a spin-mass correlation with three regimes: low spins ($a_{*}\lesssim 0.3$) for low BH masses, a steep increase with transition mass scale $\sim 50\,{\rm M}_{\odot}$ and corresponding spin $\sim 0.5$, and a saturating plateau for high masses.
For the typical Cosmic Gems cluster mass, $10^{6}\,{\rm M}_{\odot}$, and size, $1\,{\rm pc}$, the saturating spin reaches a median value $\bar{a}_{*}^{\rm sat}=0.90$ for $m_{\rm BH}\geq 100\,{\rm M}_{\odot}$, with the $1\sigma$ range covering $a_{*}=0.70-0.96$ (see Table 1). Remarkably, these BH spin values and corresponding masses are in good agreement with GW231123 (Abac et al., 2025), suggesting that some heavy BHs within the upper BH mass gap could have formed via accretion within the first compact massive stellar clusters, during their formation stage in the early Universe.
@@
Finally, we stress that the stochastic nature of the process also underlies the presence of low-spin, high-mass outliers for sufficiently massive clusters representative of the Cosmic Gems population. Specifically, we expect $\sim 1$ mass-gap BH per cluster with low spin $a_{*}\lesssim 0.1$. Massive BHs ($m_{\rm BH}\gtrsim 500\,{\rm M}_{\odot}$) with low spin $a_{*}\lesssim 0.3$ are significantly more rare but cannot be excluded.
Thus, our proposed BH spin and mass growth channel has two distinctive signatures: primarily, the spin-mass correlation of Figure 5(b), mathematically expressed by the characteristic saturating exponential profile of the BH spin distribution, as in Eq. (31); second, the presence of low-spin, high-mass BH outliers, which are difficult to explain with the repeated-merger channel. Possible important avenues for future work include (i) the analysis of GW data to identify potential candidates of our proposed channel based on these two signatures, and to distinguish them from those produced by the repeated-merger channel; (ii) the calculation of the contribution to the GW background generated by accreting BBHs in such gaseous proto-stellar clusters; (iii) the calculation of IMBH binary merger rates in these environments, which is important for the planned LISA mission.
@@ Acknowledgements.
Appendix A Stochastic kicks timestep
@@
There is a hierarchy behind Eq. (32). Dynamical friction is a cumulative relaxation process for each BH and therefore should operate on a timescale longer than the local dynamical time. It is natural then to update the stochastic forcing on the timescale over which the local perturber population evolves, $\Delta t_{\rm stoch}=\mathscr{O}(\tau_{\rm grav})$, as long as $\tau_{\rm df}\gg\tau_{\rm grav}$. In all our runs we have typically $\tau_{\rm grav}\lesssim 10^{-2}\tau_{\rm df}$, and at least $\tau_{\rm grav}<10^{-1}\tau_{\rm df}$, throughout the whole time evolution, as illustrated in Fig. 6 for two respresantive BHs of different mass. Accordingly, $\tau_{\rm grav}$ sets the stochastic timestep for $\sim 99\%$ of all kicks for BHs with final mass $m_{\bullet}\geq 15{\rm M_{\odot}}$ and for the vast majority $\gtrsim 75\%$ of BHs with lower final masses. The crossing time $\tau_{\rm cross}$ becomes relevant only for low-mass BHs displaced to the outskirts of the core after a previous strong kick, but still getting $\tau_{\rm cross}\approx\tau_{\rm grav}$ in those cases. The $\tau_{\rm turb}$ remains always larger than $\tau_{\rm grav}$ for our adopted turbulence Mach number $\mathcal{M}_{\rm turb}=2$, and more generally for any $\mathcal{M}_{\rm turb}\lesssim 10$. We also implemented in the code prescribed bounds on $\Delta t_{\rm stoch}$ to improve numerical stability, which nevertheless were never violated by Eq. (32) in practice.
@@
Figure 6: The evolution of timescales used in equation (32) for two heavy BHs grown to final masses (a) $100\,{\rm M}_{\odot}$ and (b) $985\,{\rm M}_{\odot}$, in an indicative run of our typical cluster $M_{\star}=10^{6}\,{\rm M}_{\odot}$, $r_{c,\star}=1\,{\rm pc}$, $\varepsilon=0.35$.
Figure 7: The evolution of the median stochastic timestep $\tilde{\Delta t}_{\rm stoch}$ and its $1\sigma$ percentile for final low-mass ($10-50\,{\rm M}_{\odot}$) and heavy ($50\,{\rm M}_{\odot}$) BHs. Two runs of a cluster as in Figure 6.
@@
Among heavy BHs with $m_{\bullet,{\rm f}}\geq 50\,{\rm M}_{\odot}$, the $\Delta t_{\rm stoch}$ presents negligible variation throughout the cluster evolution. Lower-mass BHs show a larger spread, because they receive larger velocity kicks and also are born at different times (originating in diverse stelar evolution channels). This is manifested as a sharp broadening of their distribution depicted in Figure 7, where we plot the evolution of the median $\tilde{\Delta t}_{\rm stoch}$ and the $1\sigma$ percentile for light and heavy BHs. Still, for all BHs, it is $\Delta t_{\rm stoch}=\mathscr{O}(10^{-2}\,{\rm Myr})$ varying less than half an order of magnitude. This is true also for all cluster parameters inspected here, with lower compactness at the same cluster mass leading to slightly higher $\Delta t_{\rm stoch}$, but still of the same order of magnitude.
Finally, we remark that heavy BHs tend to receive more kicks than lower-mass BHs, having shorter stochatstic timesteps and being born mostly earlier in cluster’s life. For our typical massive compact cluster, $M_{\star}=10^{6}\,{\rm M}_{\odot}$, $r_{c,\star}=1\,{\rm pc}$, we report the median, the range and $1\sigma$ percentile (in parenthesis) of stochastic cycles for several masses as follows: (i) the BHs which have grown massive, $(150-10^{3})\,{\rm M}_{\odot}$, have received $\approx 1400^{+180(130)}_{-200(120)}$ kicks, (ii) the BHs with final mass within the mass gap, $(50-150)\,{\rm M}_{\odot}$, have received $\approx 1300^{+290(170)}_{-190(120)}$ kicks, (iii) the low-mass BHs, $m_{\bullet,{\rm f}}<50\,{\rm M}_{\odot}$, have received $\approx 600^{+990(300)}_{-500(260)}$ kicks. The primary cluster parameter that determines the stochastic cycles is the cluster compactness rather than the cluster mass. Less compact clusters that can drive BH mass growth only up to $m_{\bullet,{\rm max}}\lesssim 150\,{\rm M}_{\odot}$ generate $\sim 25\%$ less number of kicks per BH.
@@ Appendix B Accretion disk timescales
In practice, this timescale is mostly irrelevant since the case $J_{\rm d}/(2J_{\rm BH})>1$ refers primarily to more massive BHs for which $\dot{m}>0.3$, i.e. slim disks (see equation (24)).
@@
Note that when the BH dominates the angular momentum, $J_{\rm d}/(2J_{\rm BH})<1$, it is the inner disk which reorients itself to (or counter to) the BH spin. In this case, equation (33) or (34), for the thin or slim disk respectively, should be used (instead of (36)) substituting $R$ with an inner radius $R_{\rm in}\approx R_{\rm ISCO}$. This yields $t_{\rm al,d}\ll\Delta t_{\rm stoch}$ for the parameter space of our study.

### 29 Apr 2026 v2

@@ 5 Conclusion
We calculated the BH spin distribution generated by accretion in gaseous proto-stellar clusters within $\sim 10\,{\rm Myr}$ as gas is exponentially depleted by stellar feedback. We mainly considered parameter values consistent with JWST observations of five proto-stellar clusters in the Cosmic Gems arc galaxy (Adamo et al., 2024) and also explored a broader parameter space.
Our accretion model incorporates the repeated disruption of BH accretion disks by gravitational perturbations in dense cluster environments. The disk orientation is determined by the acquired BH’s azimuthal velocity, and therefore each disruption reorients the disk. Our major results are summarized in Tables 1 and 2, Fig. 5(b), and Eq. (31).
This equation describes a BH spin distribution as a function of BH mass with a remarkably consistent functional form; it persists for all cluster parameter values investigated here, with representative ranges given in Table 1. This behavior, depicted graphically in Fig. 5(b), reveals a spin-mass correlation with three regimes: low spins ($a_{*}\lesssim 0.3$) for low BH masses, a steep increase with a transition mass scale of $\sim 50\,{\rm M}_{\odot}$ and a corresponding spin of $\sim 0.5$, and a saturating plateau for high masses.
For the typical Cosmic Gems cluster mass, $10^{6}\,{\rm M}_{\odot}$, and size, $1\,{\rm pc}$, the saturating spin reaches a median value of $\bar{a}_{*}^{\rm sat}=0.90$ for $m_{\rm BH}\geq 100\,{\rm M}_{\odot}$, with the $1\sigma$ range covering $a_{*}=0.70-0.96$ (see Table 1). Remarkably, these BH spin values and corresponding masses are in good agreement with GW231123 (Abac et al., 2025), suggesting that some heavy BHs within the upper BH mass gap could have formed via accretion within the first compact massive stellar clusters, during their formation stage in the early Universe.
@@
Finally, we stress that the stochastic nature of the process also underlies the presence of low-spin, high-mass outliers for sufficiently massive clusters representative of the Cosmic Gems population. Specifically, we expect $\sim 1$ mass-gap BH per cluster with a low spin of $a_{*}\lesssim 0.1$. Massive BHs ($m_{\rm BH}\gtrsim 500\,{\rm M}_{\odot}$) with a low spin of $a_{*}\lesssim 0.3$ are significantly rarer but cannot be excluded.
Thus, our proposed BH spin and mass growth channel has two distinctive signatures: primarily, the spin-mass correlation of Fig. 5(b), mathematically expressed by the characteristic saturating exponential profile of the BH spin distribution, as in Eq. (31); second, the presence of low-spin, high-mass BH outliers, which are difficult to explain with the repeated-merger channel. Possible important avenues for future work include (i) the analysis of GW data to identify potential candidates of our proposed channel based on these two signatures, and to distinguish them from those produced by the repeated-merger channel; (ii) the calculation of the contribution to the GW background generated by accreting BBHs in such gaseous proto-stellar clusters; and (iii) the calculation of IMBH binary merger rates in these environments, which is important for the planned LISA mission.
@@ Acknowledgements.
Appendix A Stochastic kick timesteps
@@
There is a hierarchy behind Eq. (32). Dynamical friction is a cumulative relaxation process for each BH and therefore should operate on a timescale longer than the local dynamical time. It is natural then to update the stochastic forcing on the timescale over which the local perturber population evolves, $\Delta t_{\rm stoch}=\mathscr{O}(\tau_{\rm grav})$, as long as $\tau_{\rm df}\gg\tau_{\rm grav}$. In all our runs we have typically $\tau_{\rm grav}\lesssim 10^{-2}\tau_{\rm df}$, and at least $\tau_{\rm grav}<10^{-1}\tau_{\rm df}$, throughout the whole time evolution, as illustrated in Fig. 6 for two represantive BHs of different mass. Accordingly, $\tau_{\rm grav}$ sets the stochastic timestep for $\sim 99\%$ of all kicks for BHs with final mass $m_{\bullet}\geq 15{\rm M_{\odot}}$ and for the vast majority $\gtrsim 75\%$ of BHs with lower final masses. The crossing time $\tau_{\rm cross}$ becomes relevant only for low-mass BHs displaced to the outskirts of the core after a previous strong kick, but still getting $\tau_{\rm cross}\approx\tau_{\rm grav}$ in those cases. The $\tau_{\rm turb}$ remains always larger than $\tau_{\rm grav}$ for our adopted turbulence Mach number $\mathcal{M}_{\rm turb}=2$, and more generally for any $\mathcal{M}_{\rm turb}\lesssim 10$. We also implemented in the code prescribed bounds on $\Delta t_{\rm stoch}$ to improve numerical stability, which nevertheless were never violated by Eq. (32) in practice.
@@
Figure 6: The evolution of timescales used in Eq. (32) for two heavy BHs grown to final masses (a) $100\,{\rm M}_{\odot}$ and (b) $985\,{\rm M}_{\odot}$, in an indicative run of our typical cluster $M_{\star}=10^{6}\,{\rm M}_{\odot}$, $r_{c,\star}=1\,{\rm pc}$, $\varepsilon=0.35$.
Figure 7: Evolution of the median stochastic timestep $\tilde{\Delta t}_{\rm stoch}$ and its $1\sigma$ percentile for final low-mass ($10-50\,{\rm M}_{\odot}$) and heavy ($50\,{\rm M}_{\odot}$) BHs. Two runs of a cluster as in Fig. 6.
@@
Among heavy BHs with $m_{\bullet,{\rm f}}\geq 50\,{\rm M}_{\odot}$, the $\Delta t_{\rm stoch}$ presents negligible variation throughout the cluster evolution. Lower-mass BHs show a larger spread, because they receive larger velocity kicks and also are born at different times (originating in diverse stelar evolution channels). This is manifested as a sharp broadening of their distribution depicted in Fig. 7, where we plot the evolution of the median $\tilde{\Delta t}_{\rm stoch}$ and the $1\sigma$ percentile for light and heavy BHs. Still, for all BHs, it is $\Delta t_{\rm stoch}=\mathscr{O}(10^{-2}\,{\rm Myr})$ varying less than half an order of magnitude. This is true also for all cluster parameters inspected here, with lower compactness at the same cluster mass leading to slightly higher $\Delta t_{\rm stoch}$, but still of the same order of magnitude.
Finally, we remark that heavy BHs tend to receive more kicks than lower-mass BHs, having shorter stochastic timesteps and being born mostly earlier in cluster’s life. For our typical massive compact cluster, $M_{\star}=10^{6}\,{\rm M}_{\odot}$, $r_{c,\star}=1\,{\rm pc}$, we report the median, the range and $1\sigma$ percentile (in parenthesis) of stochastic cycles for several masses as follows: (i) the BHs which have grown massive, $(150-10^{3})\,{\rm M}_{\odot}$, have received $\approx 1400^{+180(130)}_{-200(120)}$ kicks, (ii) the BHs with final mass within the mass gap, $(50-150)\,{\rm M}_{\odot}$, have received $\approx 1300^{+290(170)}_{-190(120)}$ kicks, (iii) the low-mass BHs, $m_{\bullet,{\rm f}}<50\,{\rm M}_{\odot}$, have received $\approx 600^{+990(300)}_{-500(260)}$ kicks. The primary cluster parameter that determines the stochastic cycles is the cluster compactness rather than the cluster mass. Less compact clusters that can drive BH mass growth only up to $m_{\bullet,{\rm max}}\lesssim 150\,{\rm M}_{\odot}$ generate $\sim 25\%$ less number of kicks per BH.
@@ Appendix B Accretion disk timescales
In practice, this timescale is mostly irrelevant since the case $J_{\rm d}/(2J_{\rm BH})>1$ refers primarily to more massive BHs for which $\dot{m}>0.3$, i.e. slim disks (see Eq. 24).
@@
Note that when the BH dominates the angular momentum, $J_{\rm d}/(2J_{\rm BH})<1$, it is the inner disk which reorients itself to (or counter to) the BH spin. In this case, Eqs. (33) or (34), for the thin or slim disk respectively, should be used (instead of Eq. 36) substituting $R$ with an inner radius $R_{\rm in}\approx R_{\rm ISCO}$. This yields $t_{\rm al,d}\ll\Delta t_{\rm stoch}$ for the parameter space of our study.